OpenStudy (anonymous):

Using only 2's and any combination of mathematical signs,symbols and functions can you make you 5? Brain Teaser

OpenStudy (anonymous):

That's one way but do you know another one

OpenStudy (anonymous):

@Nurali that's what i got

OpenStudy (nurali):

SQRT(.2 ^ -2) Take .2 and raise it to the power of -2 and then take the square root.

OpenStudy (nurali):

\[\sqrt{\frac{ 1 }{ 0.2^{2} }}\]

OpenStudy (nurali):

\[\sqrt{\frac{ 1 }{ 0.04 }}\]

OpenStudy (nurali):

\[\sqrt{25}\]

OpenStudy (nurali):

\[5\]

OpenStudy (nurali):

Medal?